OpenStudy (anonymous):

a,b,c are positive real numbers with abc=1 Prove: 1/(a^3+b^3+1)+1/(b^3+c^3+1)+1/(a^3+c^3+1) ≤1

OpenStudy (anonymous):

Prove \[\frac{ 1 }{ c^3+b^3+1 }+\frac{ 1 }{b^3+a^3+1 }+\frac{ 1 }{ a^3+c^3+1 }\le1\]